# CAST-LUT: Tokenizer-Guided HSV Look-Up Tables for Purple Flare Removal

### v1

@@ Introduction
In complex lighting environments, images are often affected by various color artifacts (yu2017low; abbasi2024enhancing), with the purple flare being one of the most common. Unlike general image degradations such as noise or blur, purple flare presents a unique challenge due to its diffuse nature, color specificity, and strong spatial correlation with highlight regions, severely impacting the image’s color fidelity (min2025exploring; chen2024topiq). This issue is especially prominent in mobile photography, consumer-grade cameras, and low-cost optical systems, primarily caused by factors such as internal lens reflections, sensor saturation, and chromatic dispersion (spencer1995physically).
@@
Recently, image enhancement techniques have made significant progress in areas such as deraining (liu2025deraings), dehazing (sabitha2024restoration; kumari2024real; agrawal2024fog), and color restoration (chen2021pre; liang2021swinir). However, the purple flare has not been systematically studied yet. Early attempts relied on traditional methods utilizing decoupled color spaces like HSV or Lab (lee2011purple; kim2010detection; ware2018measuring). These methods are constrained by their reliance on hand-crafted, fixed thresholds, which causes them to lack flexibility and generalization capability in complex scenes, limiting their practical application. To overcome the limitations of traditional methods, researchers have turned to deep learning. Wu et al. (wu2021train) pioneered the application of end-to-end CNNs to general flare removal by constructing a large-scale semi-synthetic dataset. To further bridge the domain gap between synthetic data and real-world flares, Deng et al. (deng2024towards) proposed a knowledge-driven hybrid approach that utilizes a flare-level estimator as prior knowledge to guide the restoration network. While these methods represent the state-of-the-art, they are essentially pixel-to-pixel restoration networks operating in the coupled RGB color space, lacking precise control over color decoupling for specific chromatic artifacts. Among them, driven by the need for efficiency, LUT-based methods have emerged as a prominent choice for lightweight architectures. However, existing LUT methods rely on the RGB space and are not sensitive to capturing the purple flare of local region transformation (see Fig. 1(a)). We observed that the HSV space can amplify the characteristic of local purple flare. Still, the non-linear transformation from RGB to HSV can amplify local color space noise, creating new artifacts in the final result (gevers2012color). To leverage the benefits of the decoupled HSV color space without incurring the instability of traditional conversion, we propose CAST-LUT. This method employs a Chroma-Aware Spectral Tokenizer (CAST) to generate stable, high-level semantic representations of the purple flare. Fig. 1(b) illustrates the critical guiding role of the CAST module. These semantic tokens then guide the generation of independent 1D-LUTs for the H, S, and V channels.
@@
Chromatic Aberration and Flare Removal. Traditional methods note limitations in processing purple flare directly in sRGB space, as intense luminance variations interfere with color assessment. Researchers have explored other color spaces (malik2018iterative). Chung et al. (chung2009detecting) identified chromatic fringes by comparing R, G, B channel intensities but faced false positives due to ambiguous purple definitions. To better separate chrominance from luminance, subsequent works shifted to other color spaces. Kim et al. (kim2010detection) used CIExy for luminance normalization, while Ju et al. leveraged YCbCr’s chrominance channels (Cb-Cr) to analyze color deviations independently (ju2013colour). However, a common bottleneck is their reliance on static, experimentally determined absolute thresholds to filter high-contrast or near-saturated areas. Given varied purple flare characteristics across cameras and scenes, fixed thresholds lack universality and robustness. Considering degraded image representation reconstruction ability, we propose using a word segmenter to aid in detail reconstruction.
Discrete Semantic Representation with Vector Quantization. Vector Quantization (VQ) is widely used to generate semantic tokens, a key component of modern perceptual systems. Its core principle maps high-dimensional feature vectors to a finite set of learned codebook embeddings. This technique gained prominence with VQ-VAE (van2017neural), which established a discrete latent space in an autoencoder architecture and integrated a discrete latent bottleneck into such frameworks. A discrete latent space offers significant advantages: it enables learning rich, semantically meaningful representations, where each codebook entry captures distinct perceptual patterns analogous to vocabulary words (schwettmann2021toward). Additionally, discretization improves generative model stability and efficiency by mitigating issues like posterior collapse, a challenge in traditional VAEs (xue2019supervised). The discretization of visual data into tokens has enabled powerful architectures like Transformers (han2022survey) and inspired token or prompt-guided approaches for other vision tasks, such as few-shot learning (li2024knn; li2025vt).
@@
LUT-based Color Image Enhancement. Recently, the LUT paradigm has shown significant potential across image restoration tasks, evolving into complex task-specific architectures. For color image denoising, where high color fidelity is critical, early LUT methods were limited by independent color channel processing. Therefore, DnLUT (yang2025dnlut) introduced a Pairwise Channel Mixer, which explicitly models inter-channel correlations by processing RGB channel pairs in parallel, boosting color denoising performance. For complex mixed-degradation scenarios (e.g., low-light enhancement), the LUT framework has advanced further. DPLUT (lin2025dplut) proposed a two-stage approach, using a Light-Adjustment LUT and Noise-Suppression LUT to decouple brightening and denoising sub-tasks. However, despite SOTA performance in their domains, these advanced LUT methods are not designed for purple flare removal. Their model priors are inconsistent with the unique characteristics and spatial distribution of flare, and they are confined to operating in the RGB space.
@@
Category Method
Venue &
Year
Global Quality
Metrics
Artifact Specific
Metrics
Efficiency
Metrics
($256^{2}$)
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ $\Delta E$ $\downarrow$ PSNR-F $\uparrow$ PSNR-NF $\uparrow$ HAE $\downarrow$
FLOPs
(G) $\downarrow$
Runtime
(ms) $\downarrow$
LUT-based 3DLUT (zeng2020learning) TPAMI’20 30.34 0.96 0.06 4.17 23.87 30.34 19.12 26.34 9.36
SR-LUT (jo2021practical) CVPR’21 29.76 0.85 0.12 5.66 21.35 29.58 10.21 25.39 8.25
SPF-LUT (li2024look) CVPR’24 32.19 0.94 0.09 5.18 21.93 33.19 7.57 34.38 9.12
NILUT (conde2024nilut) AAAI’24 32.31 0.95 0.07 4.50 24.86 32.93 4.87 30.89 7.12
DnLUT (yang2025dnlut) CVPR’25 30.52 0.88 0.10 5.17 23.31 30.25 9.51 25.16 8.71
CAST-LUT (Ours) - 34.96 0.99 0.03 2.71 30.74 34.96 4.10 23.32 6.09
Classical CBM3D (dabov2007color) TIP’07 30.57 0.90 0.22 4.73 27.49 31.57 18.53 - 43.11
MC-WNNM (xu2017multi) ICCV’17 27.98 0.80 0.20 9.53 26.23 28.00 18.01 - 121.23
Scene
Specific
Restoration
Zero-DCE (guo2020zero) CVPR’20 14.92 0.84 0.13 19.59 13.34 14.92 40.16 34.79 10.73
RUAS (liu2021retinex) CVPR’21 19.96 0.68 0.35 31.30 8.73 9.96 28.49 43.84 11.34
LightenDiffusion (Jiang_2024_ECCV) ECCV’24 30.59 0.97 0.08 5.01 24.66 30.59 27.25 27.01 10.25
BPAM (lou2025learning) ICCV’25 29.93 0.95 0.12 5.10 27.26 28.93 24.75 15.21 13.04
HVI-CIDNet (yan2025hvi) CVPR’25 32.17 0.98 0.04 2.92 26.10 33.17 17.39 27.50 14.94

### v2

@@ Introduction
In complex lighting environments, images are often affected by various color artifacts (Yu and Zhu 2017; Abbasi et al. 2024), with the purple flare being one of the most common. Unlike general image degradations such as noise or blur, purple flare presents a unique challenge due to its diffuse nature, color specificity, and strong spatial correlation with highlight regions, severely impacting the image’s color fidelity (Min et al. 2025; Chen et al. 2024). This issue is especially prominent in mobile photography, consumer-grade cameras, and low-cost optical systems, primarily caused by factors such as internal lens reflections, sensor saturation, and chromatic dispersion (Spencer et al. 1995).
@@
Recently, image enhancement techniques have made significant progress in areas such as deraining (Liu et al. 2025), dehazing (Sabitha and Eluri 2024; Kumari et al. 2024; Agrawal et al. 2024), and color restoration (Chen et al. 2021; Liang et al. 2021). However, the purple flare has not been systematically studied yet. Early attempts relied on traditional methods utilizing decoupled color spaces like HSV or Lab (Lee et al. 2011; Kim and Park 2010; Ware et al. 2018). These methods are constrained by their reliance on hand-crafted, fixed thresholds, which causes them to lack flexibility and generalization capability in complex scenes, limiting their practical application. To overcome the limitations of traditional methods, researchers have turned to deep learning. Wu et al. (2021) pioneered the application of end-to-end CNNs to general flare removal by constructing a large-scale semi-synthetic dataset. To further bridge the domain gap between synthetic data and real-world flares, Deng et al. (2024) proposed a knowledge-driven hybrid approach that utilizes a flare-level estimator as prior knowledge to guide the restoration network. While these methods represent the state-of-the-art, they are essentially pixel-to-pixel restoration networks operating in the coupled RGB color space, lacking precise control over color decoupling for specific chromatic artifacts. Among them, driven by the need for efficiency, LUT-based methods have emerged as a prominent choice for lightweight architectures. However, existing LUT methods rely on the RGB space and are not sensitive to capturing the purple flare of local region transformation (see Fig. 1(a)). We observed that the HSV space can amplify the characteristic of local purple flare. Still, the non-linear transformation from RGB to HSV can amplify local color space noise, creating new artifacts in the final result (Gevers et al. 2012). To leverage the benefits of the decoupled HSV color space without incurring the instability of traditional conversion, we propose CAST-LUT. This method employs a Chroma-Aware Spectral Tokenizer (CAST) to generate stable, high-level semantic representations of the purple flare. Fig. 1(b) illustrates the critical guiding role of the CAST module. These semantic tokens then guide the generation of independent 1D-LUTs for the H, S, and V channels.
@@
Chromatic Aberration and Flare Removal. Traditional methods note limitations in processing purple flare directly in sRGB space, as intense luminance variations interfere with color assessment. Researchers have explored other color spaces (Malik and Karthik 2018). Chung et al. (2009) identified chromatic fringes by comparing R, G, B channel intensities but faced false positives due to ambiguous purple definitions. To better separate chrominance from luminance, subsequent works shifted to other color spaces. Kim et al. (2010) used CIExy for luminance normalization, while Ju et al. leveraged YCbCr’s chrominance channels (Cb-Cr) to analyze color deviations independently (Ju and Park 2013). However, a common bottleneck is their reliance on static, experimentally determined absolute thresholds to filter high-contrast or near-saturated areas. Given varied purple flare characteristics across cameras and scenes, fixed thresholds lack universality and robustness. Considering degraded image representation reconstruction ability, we propose using a word segmenter to aid in detail reconstruction.
Discrete Semantic Representation with Vector Quantization. Vector Quantization (VQ) is widely used to generate semantic tokens, a key component of modern perceptual systems. Its core principle maps high-dimensional feature vectors to a finite set of learned codebook embeddings. This technique gained prominence with VQ-VAE (Van Den Oord et al. 2017), which established a discrete latent space in an autoencoder architecture and integrated a discrete latent bottleneck into such frameworks. A discrete latent space offers significant advantages: it enables learning rich, semantically meaningful representations, where each codebook entry captures distinct perceptual patterns analogous to vocabulary words (Schwettmann et al. 2021). Additionally, discretization improves generative model stability and efficiency by mitigating issues like posterior collapse, a challenge in traditional VAEs (Xue et al. 2019). The discretization of visual data into tokens has enabled powerful architectures like Transformers (Han et al. 2022) and inspired token or prompt-guided approaches for other vision tasks, such as few-shot learning (Li et al. 2024a, 2025).
@@
LUT-based Color Image Enhancement. Recently, the LUT paradigm has shown significant potential across image restoration tasks, evolving into complex task-specific architectures. For color image denoising, where high color fidelity is critical, early LUT methods were limited by independent color channel processing. Therefore, DnLUT (Yang et al. 2025) introduced a Pairwise Channel Mixer, which explicitly models inter-channel correlations by processing RGB channel pairs in parallel, boosting color denoising performance. For complex mixed-degradation scenarios (e.g., low-light enhancement), the LUT framework has advanced further. DPLUT (Lin et al. 2025) proposed a two-stage approach, using a Light-Adjustment LUT and Noise-Suppression LUT to decouple brightening and denoising sub-tasks. However, despite SOTA performance in their domains, these advanced LUT methods are not designed for purple flare removal. Their model priors are inconsistent with the unique characteristics and spatial distribution of flare, and they are confined to operating in the RGB space.
@@
Category Method Venue & Year Global Quality Metrics Artifact Specific Metrics Efficiency Metrics ($256^{2}$)
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ $\Delta E$ $\downarrow$ PSNR-F $\uparrow$ PSNR-NF $\uparrow$ HAE $\downarrow$ FLOPs (G) $\downarrow$ Runtime (ms) $\downarrow$
LUT-based 3DLUT (Zeng et al. 2020) TPAMI’20 30.34 0.96 0.06 4.17 23.87 30.34 19.12 26.34 9.36
SR-LUT (Jo and Kim 2021) CVPR’21 29.76 0.85 0.12 5.66 21.35 29.58 10.21 25.39 8.25
SPF-LUT (Li et al. 2024b) CVPR’24 32.19 0.94 0.09 5.18 21.93 33.19 7.57 34.38 9.12
NILUT (Conde et al. 2024) AAAI’24 32.31 0.95 0.07 4.50 24.86 32.93 4.87 30.89 7.12
DnLUT (Yang et al. 2025) CVPR’25 30.52 0.88 0.10 5.17 23.31 30.25 9.51 25.16 8.71
CAST-LUT (Ours) - 34.96 0.99 0.03 2.71 30.74 34.35 4.10 23.32 6.09
Classical CBM3D (Dabov et al. 2007) TIP’07 30.57 0.90 0.22 4.73 27.49 31.57 18.53 - 43.11
MC-WNNM (Xu et al. 2017) ICCV’17 27.98 0.80 0.20 9.53 26.23 28.00 18.01 - 121.23
Scene Specific Restoration Zero-DCE (Guo et al. 2020) CVPR’20 14.92 0.84 0.13 19.59 13.34 14.92 40.16 34.79 10.73
RUAS (Liu et al. 2021) CVPR’21 19.96 0.68 0.35 31.30 8.73 9.96 28.49 43.84 11.34
LightenDiffusion (Jiang et al. 2024) ECCV’24 30.59 0.97 0.08 5.01 24.66 30.59 27.25 27.01 10.25
BPAM (Lou et al. 2025) ICCV’25 29.93 0.95 0.12 5.10 27.26 28.93 24.75 15.21 13.04
HVI-CIDNet (Yan et al. 2025) CVPR’25 32.17 0.98 0.04 2.92 26.10 33.17 17.39 27.50 14.94